OpenStudy (anonymous):

A fish recipe suggests that for every 3/4 of an inch of thickness, the fish should be cooked for 7 1/2 minutes. If a sea bass steak is 2 1/3 inches tick, how long should it be cooked?

OpenStudy (anonymous):

This actually isn't that hard, I just got kind of confused when reading it

OpenStudy (anonymous):

And now I'm lost Can someone help please?

OpenStudy (marihelenh):

Divide 2 1/3 by 3/4 first.

OpenStudy (anonymous):

@marihelenh In fraction form or decimal?

OpenStudy (marihelenh):

Fraction will make it much easier

OpenStudy (calculusxy):

\(\large \frac{\text{thickness}}{\text{minutes}} = \frac{\text{thickness}}{\text{minutes}}\) set up a proportion

OpenStudy (anonymous):

Ermmm... I believe it's uh, somewhere near 3 and 1/ 9?

OpenStudy (marihelenh):

Yes, that is correct. That is also written 29/9.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

in improper form

OpenStudy (marihelenh):

Now multiply 3 1/9 and 7 1/2.

OpenStudy (anonymous):

I don't really see where ur going with this but... 23 and 1/3 According to the calculator

OpenStudy (anonymous):

@marihelenh ?

OpenStudy (mathstudent55):

You can think of this problem two ways. 1. use unit rate 2. use proportion

OpenStudy (calculusxy):

\[\large \frac{ \frac{ 3 }{ 4 } }{ \frac{ 15 }{ 2 } } = \frac{ \frac{ 7 }{ 3 } }{ x }\] \[\frac{ 15 }{ 2 } \times \frac{ 7 }{ 3 } = \frac{ 35 }{ 2 }\] \[\frac{ 35 }{ 2 }=\frac{ 3 }{ 4 }x \rightarrow \frac{ 35 }{ 2 } \times \frac{ 4 }{ 3 } = \frac{ 70 }{ 3} = 23\frac{ 1 }{ 3 }\]

OpenStudy (marihelenh):

Yes. That is the time. You multiplied the thickness by the time for a unit of thickness. What is 1/3 of a minute? How many seconds?

OpenStudy (calculusxy):

\(x = 23\frac{1}{3}\) i guess it's \(23\frac{1}{3} \text{minutes}\)

OpenStudy (anonymous):

Umm, how would I calculate 1/3 of a minute?

OpenStudy (marihelenh):

There are 60 seconds in a minute. 1/3 times 60 is...?

OpenStudy (mathstudent55):

1. unit rate Each 3/4 in of thickness needs 7.5 minutes. 2 1/3 minutes is how many times 3/4 inch? We divide 2 1/3 by 3/4 to find how many times it is 3/4 inches. The answer is 3 1/3. Since 2 1/3 is the same as 3 1/3 times 3/4 inch, then you need 7 1/2 minutes for each 3/4 inch, so you need 3 1/3 times 7 1/2 2. proportion \(\dfrac{3/4~inch}{7 ~1/2 ~min} = \dfrac{2~ 1/3 ~inch}{x} \) \(3/4 x = (2 1/3)(7 1/2)\) \(x = 23 ~1/3~min\)

OpenStudy (calculusxy):

look at both @mathstudent55 's and my work and you should be off fine :) we both have the same approach (proportions)

OpenStudy (calculusxy):

for these types of questions, i suggest you to use proportions as they can be very helpful :)

OpenStudy (anonymous):

Sorry I had to go. Thanks for all your help everyone!